OpenStudy (anonymous):

Use the distance formula to write an equation for all points that are 5 units from (0,0). Sketch a graph showing these points.

OpenStudy (anonymous):

We know from the Pythagorean theorem that: \[x^2+y^2=D^2\] Where \(D\) is distance. So then \[x^2+y^2=5^2=25\]

OpenStudy (ddcamp):

Distance formula: \[D = \sqrt{(y₂-y₁)² + (x₂-x₁)²}\] Let's say that the initial point (x₁, y₁) is (0,0). We're given that D=5:

OpenStudy (anonymous):

You have the equation of: \[x^2+y^2=25\] Or \[y=\pm\sqrt{25-x^2}\] Which is actually the equation of a circle with a radius of five centred at \((0,0)\)